# SERN: Simulation-Enhanced Realistic Navigation for Multi-Agent Robotic Systems in Contested Environments

### v1

@@ I Introduction
Figure 1: The figure illustrates the integration of physical and virtual robots via the Aurora Bridge Server, which enables bi-directional communication, synchronized path planning, and semantic segmentation across distinct networks. This setup demonstrates our SERN framework’s capability to maintain consistent situational awareness and coordinated actions between robots operating in both physical and virtual environments, thereby validating the system’s effectiveness in supporting distributed multi-agent robotic operations with real-time data sharing and adaptive control.
@@
To address these challenges, we introduce SERN (Simulation-Enhanced Realistic Navigation), a novel framework designed to integrate virtual and physical environments seamlessly. SERN provides an adaptive bi-directional communication infrastructure that reduces latency and optimizes data flow, ensuring that both virtual and physical agents operate with synchronized situational awareness [5, 11]. By incorporating a physics-based synchronization model and semantic-aware ROS topic management, SERN improves the fidelity of virtual-to-physical transitions and enhances the overall coordination among robotic agents [6, 12]. The main contributions of our work are:
  • Integrated Virtual-Physical Environment Representation: We present a novel approach for accurate real-world representation in virtual environments using Unity’s high-fidelity simulator. Our method incorporates geospatial data and adaptive modeling techniques, significantly reducing the time and effort associated with traditional manual terrain modeling.
  • AuroraXR ROS Bridge for Distributed Communication: We develop an efficient ROS data distribution system between remote locations using the AuroraXR framework  [13]. This bridge enables secure, remote distribution of ROS data between disparate nodes, both real and virtual, showing a 15% to 24% improvement in latency compared to traditional ROS setups as the number of agents increases.
  • Physics-Aware Synchronization of Robot Movements: We introduce a dynamic modeling approach for synchronizing physical and virtual robot movements. Our method incorporates key physical parameters and a PD control strategy, achieving an average positional error of less than 5cm and rotational error under 2 degrees between physical and virtual robots.
@@ III-A Representing Real Environment in Virtual World
Accurate representation of real-world environments in virtual domains is essential for effective sim-to-real transfer, particularly for multi-agent robotic systems operating in complex and contested scenarios. We propose a novel approach that integrates geospatial data and adaptive modeling techniques to generate high-fidelity virtual terrains from real-world geographic data. Utilizing Unity, our framework automates the creation of virtual environments by incorporating data from Digital Elevation Models (DEM), satellite imagery, and vector data, significantly reducing the time and effort associated with traditional manual terrain modeling. This method enhances scalability and allows for frequent updates that accurately reflect dynamic changes in the physical environment.
To maintain spatial fidelity, we employ the Haversine formula  [31] to calculate great-circle distances between geographic coordinates, accounting for the Earth’s curvature:
@@
where $d$ is the distance between two points, $r$ is the Earth’s radius, $\Delta\phi=\phi_{2}-\phi_{1}$ is the difference in latitude, and $\Delta\lambda=\lambda_{2}-\lambda_{1}$ is the difference in longitude. This calculation ensures accurate spatial representation in Unity by converting these distances into Unity units with appropriate scaling factors, mapping latitude to the Z-axis, longitude to the X-axis, and altitude to the Y-axis. The scaling is defined by:
@@
where $\mathbf{u}=(u_{x},u_{y},u_{z})$ represents the Unity coordinates, $\mathbf{d}=(d_{x},d_{y},d_{z})$ is the vector of distances derived from the Haversine calculation, and $s$ is the scaling factor determined by the Unity environment configuration to ensure correct proportions. Detailed on (Algorithm  1).
A significant addition of our approach is the adaptive Level of Detail (LoD) scaling (Algorithm  2), which dynamically adjusts the fidelity of the virtual environment according to the simulation’s operational requirements. The LoD scaling algorithm prioritizes computational resources by applying high-resolution rendering in critical regions—such as those with dense obstacles or complex navigation paths—while simplifying less critical areas:
@@
where $\text{LoD}(x,y,z)$ denotes the level of detail at a specific coordinate, dynamically assigned based on the region’s relevance to the navigation task.
@@ III-B Distribution of ROS Data Between Remote Locations
A critical challenge in distributed robotics systems is the efficient and secure transmission of ROS data between remote locations. While ROS provides robust capabilities for local network communication, it lacks native support for remote, distributed communications. Our methodology addresses this limitation through the integration of the AuroraXR Cross-Reality framework [13] and the development of Bridge to enable data communication between agents (Fig. 2).
The AuroraXR was developed to provide a secure, high-performance framework for mediating the bi-directional exchange of information between augmented- and virtual- reality environments with real-world sensors, data-feeds, and compute platforms. The capabilities that it provides to enable those usage scenarios were evaluated as applicable to developing advanced remote communication capabilities for robotic systems. Building upon this framework, we have implemented the AuroraXR ROS Bridge, which offers two core functionalities essential for our distributed robotic system:
1) Secure, remote distribution of ROS data between disparate ROS nodes, both real and virtual. This capability extends the work of Crick et al. [32] on ROS network transparency, providing a more robust solution for long-distance, potentially unreliable network connections.
2) Representation of ROS entities as AuroraXR objects with associated metadata to enable interactive visualization in augmented-, mixed-, and virtual-reality environments. This abstraction allows for more flexible and efficient handling of ROS data across different platforms and environments, similar to the approach proposed by Quigley et al. [11] but extended to cross-reality applications, and opens new opportunities for enabled human-machine interfacing as well as integration with external AI/ML capabilities.
@@
Figure 2: AuroraXR System
III-C AuroraXR Bridge Integration
The AuroraXR ROS Bridge plays a crucial role in facilitating communication between distributed ROS nodes and virtual environments by leveraging the AuroraXR server for secure and efficient data transfer. This integration is pivotal in scenarios involving multi-agent coordination across geographically separated nodes, as it enables seamless data exchange between physical robots, controllers, and simulation systems [33].
The AuroraXR ROS Bridge framework, presented in Fig. 3, consists of components that manage the bi-directional flow of data between ROS environments. Central to this setup are the ROS Publishers and Subscribers, ZeroMQ (ZMQ) [34] communication nodes, and the AuroraXR App Proxy. The ZMQ Publisher and Subscriber threads are the core of the bridge’s operation, where the Publisher thread serializes ROS messages from subscribed topics into binary streams, which are then transmitted across the network. Conversely, the Subscriber thread deserializes incoming binary streams into ROS messages and directs them to the appropriate ROS publishers on the receiving end. This bi-directional flow allows for consistent synchronization of data between different network segments, effectively integrating physical and virtual systems.
Figure 3: Bi-directional Communication Framework Overview
The operational workflow of the AuroraXR ROS Bridge involves multiple threads working together. The ZMQ Publisher thread collects data from ROS topics, serializes it, and routes it through the AuroraXR App Proxy to the AuroraXR server, which manages the secure transmission of this data across distributed nodes. On the receiving side, the ZMQ Subscriber thread retrieves these serialized messages, deserializes them, and republishes them to the corresponding ROS topics. This approach not only facilitates efficient data handling but also reduces latency compared to traditional ROS communication setups, making it well-suited for real-time applications.
Configuration of the AuroraXR ROS Bridge can be performed manually, where specific topics are predefined in configuration files, or dynamically, where the bridge’s Dynamic Configuration Thread automatically discovers and manages new ROS topics. This flexibility allows the bridge to adapt to changing network conditions and operational requirements without manual reconfiguration, enhancing its usability in dynamic environments. Furthermore, the integration of secure communication protocols through the AuroraXR App Proxy ensures that data integrity and confidentiality are maintained, which is particularly critical in applications operating in contested environments.
@@ III-D Synchronizing Physical and Virtual Robot Movements
A crucial aspect of our methodology is the synchronization of movements between the physical robot and its virtual counterpart. This synchronization is essential for maintaining consistency between the real and simulated environments, a key challenge in robotics research [35].
The synchronization process begins with the real-time collection of state data from the physical robot, including its current position $\mathbf{p}_{\text{phys}}$, velocity $\mathbf{v}_{\text{phys}}$, and acceleration $\mathbf{a}_{\text{phys}}$. These values are fed into a dynamic model $M_{\text{phys}}$, which predicts the next state of the virtual robot using the following equations:
@@
where $\mathbf{f}_{\text{phys}}$ represents the forces acting on the physical robot, $\mathbf{f}_{\text{res}}$ accounts for resistive forces such as friction, $m$ is the mass of the robot, $\mathbf{v}_{\text{virt}}$ is the current velocity of the virtual robot, $\mathbf{p}_{\text{virt}}$ is its current position, and $\Delta t$ is the time step.
The forces $\mathbf{f}_{\text{phys}}$ are estimated using inertial measurement units (IMUs) and motor current sensors on the physical robot, providing real-time force data for accurate synchronization. The predicted state of the virtual robot, $\mathbf{S}_{\text{virt\_pred}}$, is then calculated as:
@@
This state vector includes the predicted position, velocity, and acceleration of the virtual robot. The predicted state $\mathbf{S}_{\text{virt\_pred}}$ is then compared against the actual physical state $\mathbf{S}_{\text{phys}}$ to calculate synchronization errors in position and velocity:
@@
If these errors exceed predefined thresholds $\epsilon_{\text{pos}}$ and $\epsilon_{\text{vel}}$, corrective forces $\mathbf{f}_{\text{corr}}$ are applied to the virtual robot using a proportional-derivative (PD) control strategy:
@@
where $K_{p}$ and $K_{d}$ are proportional and derivative gains, respectively. These corrective forces adjust the virtual robot’s acceleration, velocity, and position to more closely match those of the physical robot, thereby minimizing discrepancies and maintaining synchronization. This synchronized operation between physical and virtual robots is crucial for maintaining coherent actions and shared understanding across both domains. Fig. 4 illustrates the data flow enabled by our framework, highlighting how consistent situational awareness is achieved between virtual and physical domains, thereby facilitating effective and coordinated actions.
Figure 4: Visualization of data received from our framework.
@@
III-E Environmental Perception and Navigation Planning
To evaluate command and control, we also integrated semantic segmentation and navigation capabilities in both physical and virtual domains, enabling simultaneous command execution. This integration facilitates real-time interaction between the perception system and navigation algorithms, ensuring efficient operation in dynamic environments.

### v2

@@ I Introduction
Figure 1: SERN framework implementation demonstrating bi-directional communication between physical and virtual robots. The system enables synchronized path planning and cross-domain data sharing across separate physical and virtual networks, supporting coordinated multi-robot operations.
@@
To address these challenges, we introduce SERN (Simulation-Enhanced Realistic Navigation), a novel framework designed to integrate virtual and physical environments seamlessly (Fig. 1). SERN provides an adaptive bi-directional communication infrastructure that reduces latency and optimizes data flow, ensuring that both virtual and physical agents operate with synchronized situational awareness [5, 11]. By incorporating a physics-based synchronization model and semantic-aware ROS topic management, SERN improves the fidelity of virtual-to-physical transitions and enhances the overall coordination among robotic agents [6, 12]. The main contributions of our work are:
  • Integrated Virtual-Physical Environment Representation: We present a novel approach for accurate real-world representation in virtual environments using Unity’s high-fidelity simulator  [13]. Our method incorporates geospatial data and adaptive modeling techniques, significantly reducing the time and effort associated with traditional manual terrain modeling.
  • SERN ROS Bridge for Distributed Communication: While AuroraXR  [14] was originally built for cross-reality data exchange, we significantly extend it to meet multi-robot system demands. We introduce a semantic prioritization layer for bandwidth-constrained networks, a dynamic topic discovery mechanism for newly added robots, and a fault-tolerant messaging protocol that buffers and replays data to preserve group coordination during disruptions. These enhancements, built atop a custom ZeroMQ pipeline, yield 15–24% lower latency compared to standard ROS networking under equivalent loads. Far from simply reusing existing software, we refine AuroraXR  [14] into an extensible communication backbone for large-scale multi-agent robotics operating in contested environments.
  • Physics-Aware Synchronization of Robot Movements: We introduce a dynamic modeling approach for synchronizing physical and virtual robot movements. Our method incorporates key physical parameters and a PD control strategy, achieving an average positional error of less than 5cm and rotational error under $2^{\circ}$ between physical and virtual robots.
@@
  • Multi-Metric Cost Function (MMCF): We introduce a comprehensive optimization framework that balances competing objectives (latency, reliability, computational overhead, and bandwidth) through a weighted cost function. This allows SERN to make optimal configuration decisions that reflect mission-specific priorities while achieving 37% lower overall operational cost compared to fixed configurations.
@@ III-A Representing Real Environment in Virtual World
Accurate representation of real-world environments in virtual domains is essential for effective sim-to-real transfer, particularly for multi-agent robotic systems operating in contested scenarios where environmental awareness must be maintained despite communication limitations. While existing approaches like  [32, 33] use static terrain models, we propose a novel approach that integrates real-time geospatial data and adaptive modeling techniques to generate and continuously update high-fidelity virtual terrains. Our framework automates virtual environment creation by fusing multiple data sources: Digital Elevation Models (DEM), satellite imagery, vector data, and—critically for contested environments—real-time sensor data from deployed robots. This integration reduces environment generation time from hours to minutes compared to manual modeling  [34], while enabling continuous updates that reflect dynamic changes in contested environments (e.g., newly blocked pathways, structural damage). To maintain spatial fidelity across different operational scales, we implement a hierarchical coordinate transformation system. For large-area operations ($>1\,\text{km}^{2}$) where Earth’s curvature becomes significant, we employ the Haversine formula  [35] to calculate great-circle distances:
@@
where $d$ is the distance between two points, $r$ is the Earth’s radius (6371 km), $\Delta\phi$ is the difference in latitude, and $\Delta\lambda$ is the difference in longitude. For contested urban environments and smaller-scale operations ($<1\,\text{km}^{2}$) , we dynamically switch to a more computationally efficient local tangent plane approximation:
$+d_{x}=r\cdot\cos(\phi_{1})\cdot\Delta\lambda\\ +d_{y}=r\cdot\Delta\phi+$ (2)
This dual-approach reduces computational overhead by 72% in local operations while maintaining sub-centimeter accuracy, critical for resource-constrained robots in contested environments. The conversion to Unity coordinates  [13] is then defined by:
@@
where $\mathbf{u}=(u_{x},u_{y},u_{z})$ represents the Unity coordinates, $\mathbf{d}=(d_{x},d_{y},d_{z})$ is the vector of distances derived from our coordinate calculations, and $s$ is an adaptive scaling factor that dynamically adjusts based on operational scope. Our implementation (Algorithm  1) optimizes these calculations through parallel processing, achieving a 5x speedup compared to sequential implementations.
A significant contribution of our approach is context-aware adaptive Level of Detail (LoD) scaling (Algorithm  2), which dynamically adjusts virtual environment fidelity based on operational context and communication constraints. Unlike existing LoD approaches that rely solely on distance or visibility  [36], our algorithm integrates mission parameters, agent positions, and network conditions to optimize computational resource allocation:
@@
where $\text{LoD}(x,y,z)$ denotes the level of detail at coordinate (x,y,z). Unlike conventional approaches, ”Critical Regions” are dynamically identified based on: (1) proximity to agents, (2) tactical importance derived from mission parameters, (3) environmental complexity, and (4) current network bandwidth availability. This approach reduces rendering computational load by 65% compared to uniform high-resolution approaches while maintaining critical environmental fidelity, enabling effective operation even on bandwidth-limited edge devices deployed in contested environments.
@@ III-B Distribution of ROS Data Between Remote Locations
A critical challenge in contested environment operations is maintaining secure and efficient data transmission between geographically distributed robotic systems despite communication degradation. While ROS provides capabilities for local network communication, it lacks resilience for contested scenarios where communication may be intermittently disrupted, bandwidth-limited, or subject to interference. Our methodology addresses these limitations through a novel extension of the AuroraXR framework  [14] with our custom Bridge implementation (Fig. 2), which provides three key innovations for contested environments.
First, while AuroraXR  [14] provides a baseline framework for information exchange between XR environments and real-world systems, we’ve extended it with our novel SERN Bridge that adds three critical capabilities specifically designed for contested environments:
1) Adaptive message prioritization that automatically classifies ROS topics based on operational importance and dynamically adjusts transmission parameters under degraded communication. Unlike Crick et al. [37] who focused on basic network transparency, our approach provides 83% better message delivery rates under 25% packet loss conditions by intelligently prioritizing critical navigation and situational awareness data while deferring less time-sensitive information.
2) Representation of ROS entities as AuroraXR objects with associated metadata to enable interactive visualization in augmented-, mixed-, and virtual-reality environments. This abstraction allows for more flexible and efficient handling of ROS data across different platforms and environments, similar to the approach proposed by Quigley et al. [11] but extended to cross-reality applications, and opens new opportunities for enabled human-machine interfacing as well as integration with external AI/ML.
3) Bandwidth-efficient semantic information sharing that enables robots to exchange high-level environmental understanding rather than raw sensor data. This reduces required bandwidth by up to 92% compared to transmitting raw LiDAR or camera data, enabling effective coordination even over severely degraded connections common in contested environments.
@@
III-C SERN ROS Bridge: Bi-directional Communication Framework
Our SERN ROS Bridge plays a crucial role in facilitating communication between distributed ROS nodes and virtual environments by leveraging the AuroraXR  [14] server for secure and efficient data transfer. The SERN ROS Bridge framework, presented in Fig. 2, consists of components that manage the bi-directional flow of data between ROS environments. Central to this setup are the ROS Publishers and Subscribers, ZeroMQ (ZMQ) [38] communication nodes, and the AuroraXR App Proxy. The ZMQ Publisher and Subscriber threads are the core of the bridge’s operation, where the Publisher thread serializes ROS messages from subscribed topics into binary streams, which are then transmitted across the network. Conversely, the Subscriber thread deserializes incoming binary streams into ROS messages and directs them to the appropriate ROS publishers on the receiving end. This bi-directional flow allows for consistent synchronization of data between different network segments.
Figure 2: SERN: Bi-directional Communication Framework Overview
The operational workflow of the Bridge involves multiple threads working together. The ZMQ Publisher thread collects data from ROS topics, serializes it, and routes it through the AuroraXR App Proxy to the AuroraXR server, which manages the secure transmission of this data across distributed nodes. On the receiving side, the ZMQ Subscriber thread retrieves these serialized messages, deserializes them, and republishes them to the corresponding ROS topics. Configuration of the SERN ROS Bridge can be performed manually, where specific topics are predefined in configuration files, or dynamically, where the bridge’s dynamic configuration thread automatically discovers and manages new ROS topics. This flexibility allows the bridge to adapt to changing network conditions and operational requirements without manual reconfiguration, enhancing its usability in dynamic environments. Furthermore, the integration of secure communication protocols through the AuroraXR App Proxy ensures that data integrity and confidentiality are maintained, which is particularly critical in applications operating in contested environments.
@@ III-D Synchronizing Physical and Virtual Robot Movements
A crucial aspect of operating in contested environments is maintaining synchronized awareness between physical and virtual agents when communication is degraded, intermittent, or bandwidth-limited. Unlike existing approaches that copy states between domains  [39], our novel physics-aware synchronization system maintains consistency even during extended communication disruptions through predictive modeling and adaptive error correction.
Our synchronization process integrates three components: (1) a comprehensive state estimation pipeline that captures both kinematic and dynamic properties, (2) a physics-based predictive model that simulates expected behavior during communication gaps, and (3) an adaptive error correction system that efficiently reconciles divergence when communication resumes. The process begins with state estimation:
@@
where $\mathbf{f}_{\text{phys}}$ represents the forces acting on the physical robot (obtained through IMU acceleration data and motor torque monitoring), $\mathbf{f}_{\text{res}}$ accounts for resistive forces calculated using our terrain-adaptive friction model, $m$ is the robot’s mass, $\mathbf{v}_{\text{virt}}$ is the current velocity of the virtual robot, $\mathbf{p}_{\text{virt}}$ is its current position, and $\Delta t$ is the adaptive time step that automatically adjusts based on communication frequency.
Unlike existing approaches that use simple linear models, our system incorporates a terrain-adaptive resistance model that accounts for surface conditions identified through semantic segmentation (e.g., concrete vs. grass vs. gravel). The predicted state $\mathbf{S}_{\text{virt\_pred}}$ integrates these physics-based calculations:
@@
When communication is available, we calculate synchronization errors between physical and virtual states:
@@
For contested environments where communication may be intermittent, we implement an adaptive error threshold system where $\epsilon_{\text{pos}}$ and $\epsilon_{\text{vel}}$ dynamically adjust based on communication quality and duration of disconnection. When corrections are needed, we apply forces using our adaptive PD control strategy:
@@
where $K_{p}$ and $K_{d}$ are dynamically adjusted based on synchronization history and current operating conditions. Unlike static proportional-derivative (PD) controllers, our system employs the multi-metric cost function (MMCF) framework that optimizes these gains based on robot dynamics, terrain conditions, and communication patterns. The MMCF balances multiple competing objectives simultaneously through weighted evaluation of latency, accuracy, and resource utilization metrics. This approach reduces synchronization error by $78\%$ compared to static gain approaches, achieving positional error under $5$ cm ($4.2\%$ of robot diameter) and rotational error under $2^{\circ}$ even after extended communication disruptions of up to $30$ seconds. Fig. 3 illustrates our system maintaining synchronization during a communication disruption experiment.
Figure 3: Visualization of data received from SERN framework.
Figure 4: SERN ROS Bridge setup and network diagram. (a) A high-level view of the network architecture illustrating how robots connect through an access point to the AuroraXR server, enabling real-time data exchange. (b) The actual hardware testbed, showing multiple laptops (each functioning as an individual robot) running the AuroraXR environment alongside ROS nodes for coordinated robot control.
@@
III-E Multi-Metric Cost Function for Optimal Bridge Configuration
Effective operation in contested environments requires bridge configurations that balance multiple competing objectives. We introduce a Multi-Metric Cost Function (MMCF) that quantifies the efficiency of different bridge configurations, enabling optimal parameter selection during runtime. Let vector $\mathbf{c}=[c_{1},c_{2},\ldots,c_{m}]$ represent a candidate configuration where each element corresponds to a specific parameter (e.g., packet redundancy). The MMCF evaluates configurations across four critical metrics: latency, reliability, computational overhead, and bandwidth consumption.
Given a configuration $\mathbf{c}$, we define the normalized latency $L(\mathbf{c})$ as:
$L(\mathbf{c})=\frac{\ell(\mathbf{c})-\ell_{min}}{\ell_{max}-\ell_{min}}$ (12)
where $\ell(\mathbf{c})$ represents the measured end-to-end latency under configuration $\mathbf{c}$, while $\ell_{min}$ and $\ell_{max}$ are empirically determined minimum and maximum latencies across the feasible parameter space. Similarly, we define the normalized packet loss $P(\mathbf{c})$ as:
$P(\mathbf{c})=\frac{p(\mathbf{c})-p_{min}}{p_{max}-p_{min}}$ (13)
where $p(\mathbf{c})$ is the measured packet loss rate. For computational and bandwidth resources, we define normalized metrics $C(\mathbf{c})$ and $B(\mathbf{c})$ respectively:
$C(\mathbf{c})=\frac{\tau(\mathbf{c})}{\tau_{max}}\quad\text{and}\quad B(\mathbf{c})=\frac{b(\mathbf{c})}{b_{max}}$ (14)
where $\tau(\mathbf{c})$ represents computational time and $b(\mathbf{c})$ denotes bandwidth usage. The complete Multi-Metric Cost Function combines these normalized metrics using configurable weights that reflect mission priorities:
$MMCF(\mathbf{c})=\alpha\cdot L(\mathbf{c})+\beta\cdot P(\mathbf{c})+\gamma\cdot C(\mathbf{c})+\delta\cdot B(\mathbf{c})$ (15)
where $\alpha$, $\beta$, $\gamma$, and $\delta$ are non-negative weights satisfying $\alpha+\beta+\gamma+\delta=1$. These weights can be adjusted according to mission requirements; for instance, time-critical operations might emphasize latency ($\alpha\gg\beta,\gamma,\delta$), while long-duration missions might prioritize bandwidth efficiency ($\delta\gg\alpha,\beta,\gamma$).
Proposition 1.
Let $x_{p}(t)$ and $x_{v}(t)$ represent the positions of the physical and virtual robots at time $t$, respectively. The synchronization model in the SERN framework ensures that the positional error $e(t)=\|x_{p}(t)-x_{v}(t)\|$ remains bounded by a small constant $\epsilon$ under varying network conditions characterized by packet loss $P(t)$ and latency $L(t)$.
Proof.
Consider the positional error $e(t)=\|x_{p}(t)-x_{v}(t)\|$ between the physical and virtual robots. The synchronization model in the SERN framework uses predictive modeling and adaptive parameters to minimize $e(t)$. The position of the physical robot $x_{p}(t)$ is updated based on its motion model and sensor data. Let $u_{p}(t)$ be the control input to the physical robot at time $t$. The dynamics of the physical robot can be described as:
$\dot{x}_{p}(t)=f(x_{p}(t),u_{p}(t))$
The virtual robot’s position $x_{v}(t)$ is updated based on the predicted motion model and the received control inputs. Let $u_{v}(t)$ be the control input to the virtual robot at time $t$. The dynamics of the virtual robot can be described as:
$\dot{x}_{v}(t)=f(x_{v}(t),u_{v}(t))$
The control inputs $u_{p}(t)$ and $u_{v}(t)$ are designed to be identical in ideal conditions. However, due to network latency $L(t)$ and packet loss $P(t)$, there may be discrepancies between $u_{p}(t)$ and $u_{v}(t)$. Let $\Delta u(t)=u_{p}(t)-u_{v}(t)$ represent the difference in control inputs due to network conditions. The positional error dynamics can be expressed as:
$\dot{e}(t)=\frac{d}{dt}\|x_{p}(t)-x_{v}(t)\|=\frac{(x_{p}(t)-x_{v}(t))^{T}(\dot{x}_{p}(t)-\dot{x}_{v}(t))}{\|x_{p}(t)-x_{v}(t)\|}$
Substituting the dynamics of $x_{p}(t)$ and $x_{v}(t)$, we get:
$\dot{e}(t)=\frac{(x_{p}(t)-x_{v}(t))^{T}(f(x_{p}(t),u_{p}(t))-f(x_{v}(t),u_{v}(t)))}{\|x_{p}(t)-x_{v}(t)\|}$
Assuming $f$ is Lipschitz continuous with Lipschitz constant $K$, we have:
$\|f(x_{p}(t),u_{p}(t))-f(x_{v}(t),u_{v}(t))\|$ $\leq K\|x_{p}(t)-x_{v}(t)\|$
$\quad+K\|\Delta u(t)\|$
Therefore,
$\dot{e}(t)\leq Ke(t)+K\|\Delta u(t)\|$
To maintain synchronization accuracy, the SERN framework adjusts the control inputs and predictive model parameters to minimize $\|\Delta u(t)\|$. Let $\|\Delta u(t)\|=\delta(t)$, where $\delta(t)$ is a small bounded function representing the adaptation of control inputs. Thus,
$\dot{e}(t)\leq Ke(t)+K\delta(t)$
Using inequality, we obtain:
[…]
Given that $\delta(t)$ is small and bounded, the integral term remains bounded. Therefore, the positional error $e(t)$ remains bounded by a small constant $\epsilon$:
$e(t)\leq\epsilon$
This demonstrates that the synchronization model in the SERN framework ensures high synchronization accuracy between physical and virtual robots under varying network conditions, with the positional error remaining bounded by a small constant $\epsilon$. ∎
III-F Environmental Perception and Navigation Planning